For a given computer system, the main memory is 256Mbyte; word size is 4 bytes; block size is 64 bytes; cache size is 64 Kbytes. what is the number of cache line?

Question 3 options:

A

64Kbyte/4bytes

B

256Mbyte/4bytes

C

256Mbyte/64bytes

D

64Kbyte/64bytes

Question 4 (3 points)

Consider a magnetic disk drive with 8 double sided platters, 2000 tracks per disk surface. Each track has a capacity 2048 KBytes. Sector size is 2KBytes. What is the capacity of acylinder?

Question 4 options:

A

2000*2048 KB

B

8*2*2048 KB

C

8*2*2000*2048 KB

D

2000*2048*2 KB
Answer #1

Question 3)

Given data:

Main memory = 256MB
Word size = 4B
Block size = 64B
Cache size = 64KB

Formula for no.of cache lines:

No.of cache lines = Cache size/Block size

No.of cache lines = 64KB/64B

Option D is correct.

-------------------------------------------------------------------------------------------------------

Question 4)

Given data:

No.of double sided platters=8
No.of surfaces = No.of double sided platters*2 = 8*2 =16.
No.of track per surface= 2000.
Each track capacity = 2048KB
Sector size=2KB

Formula for cylinder capacity :

Capacity = No.of surfaces*(No.of tracks per surface)*capacity of track

Capacity = 8*2*2000*2048KB

Option C is correct.